### v1

# SMURF: Continuous Dynamics for Motion-Deblurring Radiance Fields

Abstract
Neural radiance fields (NeRF) has attracted considerable attention for their exceptional ability in synthesizing novel views with high fidelity. However, the presence of motion blur, resulting from slight camera movements during extended shutter exposures, poses a significant challenge, potentially compromising the quality of the reconstructed 3D scenes. While recent studies have addressed this issue, they do not consider the continuous dynamics of camera movements during image acquisition, leading to inaccurate scene reconstruction. Additionally, these methods are plagued by slow training and rendering speed. To effectively handle these issues, we propose sequential motion understanding radiance fields (SMURF), a novel approach that employs neural ordinary differential equation (Neural-ODE) to model continuous camera motion and leverages the explicit volumetric representation method for faster training and robustness to motion-blurred input images. The core idea of the SMURF is continuous motion blurring kernel (CMBK), a unique module designed to model a continuous camera movements for processing blurry inputs. Our model, rigorously evaluated against benchmark datasets, demonstrates state-of-the-art performance both quantitatively and qualitatively. Source code will be publicly available at https://github.com/Jho-Yonsei/SMURF.
Keywords:
Neural Rendering View Synthesis Motion Deblurring
1 Introduction
Reconstructing complex 3D scenes from 2D images of different views and re-rendering novel view images represent a core problem in computer vision and graphics, with significant applications in augmented reality (AR) and virtual reality (VR). Over the past few years, photo-realistic novel view synthesis has greatly advanced with the emergence of Neural Radiance Fields (NeRF) [33]. It takes 3D spatial coordinates and 2D viewing directions (i.e. 5D vector) as inputs for mapping to the radiance and volume density, where the process is parameterized through an implicit multi-layer perceptron (MLP) based model. Exploiting classical volume rendering techniques [19], NeRF integrates the output radiance and volume density along the emitted rays, making the volume rendering process fully differentiable.
Generally, NeRF variants have reconstructed 3D scenes using well-captured, noise-free images as inputs along with calibrated camera parameters. However, images captured in real-world may include various types of noise (e.g. camera motion blur, finite depth-of-field), which exert negative impacts on the network to train the complex geometry of the scene. Recently, many works [31, 25, 56, 40, 26] have been conducted on NeRF that takes camera motion-blurred images as input, which is caused by camera movement during the exposure time. Deblur-NeRF [31] first proposes a method that models blurring kernel by imitating the deconvolution method for blind image deblurring, to reconstruct 3D scenes from motion-blurred images and render sharp novel view images. Since the inception of Deblur-NeRF, various methods [25, 56, 40, 26] for estimating the blurring kernel have been proposed, but they have following limitations: a) their blurring kernel does not continuously model the camera movement during exposure time as shwon in Fig. 1. Since actual camera movement is normally continuous, it can be represented as a non-linear continuous function over time, and such continuous modeling allows for a more precise tracking of the camera movement path, even when the movement is complex or irregular. Previous works all use a method where the warped rays of the blurring kernel are computed jointly, making the kernel for images with complex camera motion imprecise. b) None of them overcome the tradeoff between performance, and the speed of training and rendering. DP-NeRF [25] shows performance improvement but suffers from slow training and rendering, whereas PDRF [40] shows fast training but offers relatively lower performance as shown in Fig. 2.
Figure 1: Comparison of joint kernel generation and our sequential kernel generation. A blurry image $I_{blur}$ is acquired as the camera moves over the exposure time ($t_{0}\sim t_{\mathcal{N}}$), with images $I_{t}$ captured at each camera pose being composited together. Previous methods generate warped rays of the blurring kernel in a single step by passing through a parameterized network $h_{p}$ without considering the temporal sequence of camera motion. However, our approach iteratively estimates warped rays along the sequential camera motion trajectory.
In this paper, we propose a sequential motion understanding radiance fields (SMURF), which incorporates a novel continuous motion blur kernel (CMBK) for modeling precise camera movements from blurry images. The CMBK estimates the small change in pose regarding camera motion through the view information of each scene and the 2D pixel location of the input ray. The output values of the CMBK are not computed jointly as in previous studies but are designed to be computed sequentially according to camera motion, as shown in Fig. 1. Particularly, recognizing that the sequential camera movements share a single continuous function, we employ a neural ordinary differential equation (Neural-ODE) [7] to ensure that the sequentially computed camera movements exhibits continuity. Specifically, to reflect the physics inherent in camera motion, we apply continuous dynamics in the latent space to CMBK, and transform the latent feature into changes in ray within the physical space. Sequentially estimating the small pose change with a light shared function implies that the CMBK does not overly depend on the learning parameters. Additionally, we introduce two regularization strategies to prevent the divergence of camera motion generated through CMBK. The first is residual momentum, which ensures that the warped ray does not deviate significantly from the previous ray, thereby preventing overfitting. The second is the output suppression loss, which also serves as a regularization method to ensure that the warped ray does not diverge significantly from the initial ray.
Figure 2: Comparison of performance on synthetic scene datasets in terms of training time. The performance comparison metrics are PSNR, SSIM, and LPIPS.
Our SMURF leverages explicit volumetric representation methods [18, 6, 34, 50] as its backbone. Specifically, the tensor factorization-based approach, Tensorial Radiance Fields (TensoRF) [6], facilitates compact and efficient 3D scene reconstruction. Considering the fact that some parts may exhibit significant blur while others are almost free of blur in a single image, incomplete blurred information and complete sharp information merge into a few adjacent voxels via blurring kernel and ensure a coherent 3D scene. Therefore, we adopt the TensoRF as our backbone as it effectively mitigates the uncertainty of information caused by motion blur and enables faster training and high quality rendering compared to previous approaches. The outstanding performance of SMURF relative to its fast speed is demonstrated in Fig. 2. While there exists a deblurring approach [26] that uses voxel-based radiance fields as a backbone and shares the common goal of fast training with ours, a key difference lies in their use of Plenoxels [59] as the backbone. Furthermore, while they adopt the different explicit representation method from ours with another goal of efficient memory consumption, we diverge in our objectives, as we aim to leverage the advantages of a 3D voxel-based method when utilizing blurry images as input.
To demonstrate the effectiveness of the proposed SMURF both quantitatively and qualitatively, we conduct extensive experiments on synthetic and real-world scenes. Our experimental results elucidate the advantages of the CMBK in comparison to previously presented blurring kernels.
Our main contributions are summarized as follows:
  • $\bullet$
    We propose a continuous motion blur kernel (CMBK) to sequentially estimate precise, continuous camera motion from blurry images.
  • $\bullet$
    We propose two regularization strategies that guide the warped rays to prevent divergence: residual momentum, and output suppression loss.
  • $\bullet$
    We significantly alleviate the tradeoff between performance and training speed exploiting CMBK and 3D tensor factorization-based representation.
  • $\bullet$
    Our sequential motion understanding radiance fields (SMURF) exhibit higher visual quality and quantitative performance compared to existing approaches.
2 Related Work
2.1 Neural Radiance Fields (NeRF)
The synthesis of photo-realistic novel views from images of different viewpoints has attracted considerable attention with the advancement of NeRF [33]. In particular, NeRF uses 3D spatial coordinates and 2D viewing directions to map radiance and volume density, a process facilitated by implicit neural representation (INR). Following the success of NeRF, various sub-domains considering real-world scenarios have been explored. Many recent studies have extensively applied NeRF across various fields, including dynamic 3D scene modeling [37, 28, 43, 54, 38, 27, 30, 61, 3], scene relighting [48, 2, 32, 41, 53], and 3D reconstruction [55, 51, 57]. Furthermore, recent studies have applied NeRF for deblurring 3D scenes [31, 25, 40, 56], assuming blurry input images from real-world conditions and aiming to produce clean images. Additionally, due to the slow rendering caused by the costly MLP layers required to evaluate each pixel’s density and color, some studies [23, 42, 36, 6, 34, 59, 50] design networks for faster training and rendering. To address these challenges, various voxel-based explicit rendering methods [6, 34, 59, 50] have been proposed, offering competitive performance to NeRF while significantly reducing training and rendering times. We adopt an explicit representation method, TensoRF [6] as the backbone, which facilitates fast training and achieves accurate and efficient novel-view rendering.
2.2 NeRFs for Blurry Images
Recently, several studies have been conducted to deblur 3D scenes and synthesize clean novel view images using blurry input images from different views. Deblur-NeRF [31] proposes, for the first time, an implicit blurring kernel in 3D space inspired by blind deblurring methods for 2D vision [22, 17, 52, 35, 10]. This blurring kernel is intentionally trained close to the pixels of the blurry image, and during rendering, a clean image is obtained without the trained kernel. DP-NeRF [25] models a rigid kernel that assumes physical scene priors when blurry images are captured through a camera, utilizing only the scene’s view information regardless of pixel location. BAD-NeRF [56], assuming a very short camera exposure time, designs a kernel that linearly interpolates the camera motion trajectory in SE(3) space. However, these approaches estimate the change in camera pose jointly, not modeling the fact that light traces left by camera motion during exposure time appear sequentially. Therefore, we aim to design a kernel where the elements of the blurring kernel, namely the small changes in camera pose, are estimated sequentially over time.
2.3 Neural Ordinary Differential Equations (Neural ODEs)
Neural-ODEs [7] are proposed to interpret neural networks within the framework of ordinary differential equations, representing the underlying dynamics inherent in hidden parameters. Neural-ODEs model a parameterized time-continuous latent state and output a unique solution of the integral of continuous dynamics, utilizing given initial values and various numerical differential equation solvers (e.g. Euler’s method [12], Runge-Kutta method [24], Dormand-Prince-Shampine method [11]). They are extensively used to continuously connect the latent space embedded in videos [20, 39, 13, 16] or model the continuous dynamics of irregularly sampled time-series data [44]. Inspired by the fact that camera motion is continuous, we exploit Neural-ODEs to create a precise blurring kernel that models the hidden continuous dynamics inherent in camera motion.
3 Method
In this section, we detail the optimization process of SMURF. In Sec. 3.1, we explain the preliminaries about the 3D tensor factorization-based rendering method and 3D scene blind deblurring for our methodology. Sec. 3.2 explains how continuous dynamics are applied to our CMBK to generate warped rays. Sec. 3.3 discusses the objective function and the optimization process. The framework of SMURF is illustrated in Fig. 3.
3.1 Preliminary
Tensorial Radiance Fields (TensoRF).
We follow the architecture of TensoRF [6], an explicit voxel-based volumetric representation utilizing CANDECOMP/PARAFAC (CP) [4, 14] and block term decomposition [9], which models an efficient view-dependent sparse voxel grid. TensoRF optimizes two 3D grid tensors, $\mathcal{G}_{\sigma},\mathcal{G}_{\mathbf{c}}\in\mathbb{R}^{FXYZ}$, for estimating volume density and view-dependent appearance feature, employing the vector-matrix (VM) decomposition that effectively extends CP decomposition. Exploiting this method, the grid (3rd-order) tensor is decomposed into low-rank 1D vectors and 2D matrices across three modes, reducing the space complexity from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$ compared to conventional explicit representations [59]. This grid is represented as follows:
$\mathcal{G}=\sum_{r=1}^{R_{1}}\mathbf{b}_{r}^{F_{1}}\circ\mathbf{v}_{r}^{X}\circ\mathbf{M}_{r}^{YZ}+\sum_{r=1}^{R_{2}}\mathbf{b}_{r}^{F_{2}}\circ\mathbf{v}_{r}^{Y}\circ\mathbf{M}_{r}^{XZ}+\sum_{r=1}^{R_{3}}\mathbf{b}_{r}^{F_{3}}\circ\mathbf{v}_{r}^{Z}\circ\mathbf{M}_{r}^{XY},$ (1)
where $\circ$ denotes the outer product; $R_{1},R_{2}$, and $R_{3}$ indicates the number of low-rank components. $\mathbf{b}^{F}\in\mathbb{R}^{F}$ and $\mathbf{v}^{X}\in\mathbb{R}^{X}$are the vectors along the $F$ and $X$ axes in each mode, and $\mathbf{M}^{YZ}\in\mathbb{R}^{YZ}$ is the matrix in the $XY$ plane for each mode. Once the grids are defined, the radiance field at a 3D point $\mathbf{x}$ for computing volume density $\sigma$ and color $\mathbf{c}$ is defined as follows:
$\sigma(\mathbf{x}),\mathbf{c}(\mathbf{x})=\mathcal{G}_{\sigma}(\mathbf{x}),\mathcal{S}(\mathcal{G}_{\mathbf{c}}(\mathbf{x}),d),$ (2)
where $\mathcal{S}$ denotes a parameterized shallow MLP that converts viewing direction $d$ and appearance feature $\mathcal{G}_{\mathbf{c}}(\mathbf{x})$ into color. The features obtained from the grid, $\mathcal{G}_{\sigma}(\mathbf{x})$ and $\mathcal{G}_{\mathbf{c}}(\mathbf{x})$, are trilinearly interpolated from adjacent voxels.
To render an image for a given view, TensoRF uses a differentiable classic volume rendering technique [19] with $\sigma$ and $\mathbf{c}$. The color $\mathbf{c}_{p}$ for each pixel reached by ray $\mathbf{r}$ is computed as follows:
$\mathbf{c}_{p}(\mathbf{r})=\sum_{i=1}^{N}T_{i}\left(1-\exp\left(-\sigma_{i}\delta_{i}\right)\right)\mathbf{c}_{i},~{}T_{i}=\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),$ (3)
where $T_{i}$ is the transmittance; $N$ and $\delta$ denote the number of sampled points and the step size between adjacent samples on ray $\mathbf{r}$, respectively.
Figure 3: Method overview of the SMURF. The CMBK encoder $\mathcal{E}_{\theta}$ transforms the embedded 2D pixel location $\mathbf{p}$ of the initial ray and view index into a latent feature. This feature is extended into an IVP with a parameterized derivative function $f_{\phi}$ in the latent space. Then, it is sequentially solved using a Neural-ODE along with given time $t$ and a chrono-view embedding function $\mathbf{\Psi}$, obtaining latent features for all warped rays. These features are transformed into changes of the ray (i.e., $\Delta\mathbf{p}$ and $\Delta\mathbf{o}$) through a decoder $\mathcal{D}_{\xi}$, and we get the warped rays by applying Eq. 11 to the initial ray. These rays are rendered into 2D pixel colors through a 3D grid-based method, and a blurry color is acquired by summing up the colors with weights from the decoder.
Blind Deblurring for 3D Scene.
For image blind deblurring [58, 5, 49], the blurring kernel $h$ is estimated in a fully unsupervised manner, and this process is achieved by convolving $h$ with sharp images. Ma et al. [31] apply this algorithm to NeRF, modeling an adaptive kernel $h_{p}$ for each ray,
$\mathbf{c}_{blur}(\mathbf{r})=\mathbf{c}_{p}(\mathbf{r})*h_{p}(\mathbf{r}),$ (4)
where $\mathbf{c}_{blur}$ and $\mathbf{c}_{p}$ respectively denote the color of the blurry pixel and the sharp pixel; *** indicates the convolution operator.
The conventional image deblurring kernel takes a fixed grid of size $K\times K$ centered around the location $p$. However, for 3D scene deblurring, applying such a kernel based on NeRF requires computing $\mathbf{c}_{p}$ for $K^{2}$ rays, which is inefficient in terms of memory and training time. To optimize a kernel for the 3D scene, it is necessary to produce a sparse kernel with fewer rays. Ma et al. [31] designs the sprase kernel to acquire blurry color, and for the temporally continuous camera motion, we extend the process as follows:
$\mathbf{c}_{blur}(\mathbf{r})=\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}\mathbf{c}_{p}(\mathbf{r}^{{t}_{i}}),~{}\text{w.r.t.}~{}\sum_{i=0}^{\mathcal{N}}w_{p}^{t_{i}}=1,$ (5)
where $\mathcal{N}$ and $t_{i}$ respectively denote the number of warped rays in camera motion and the instantaneous time during exposure; $w_{p}$ is the corresponding weight at each ray’s location, and $\mathbf{r}^{t_{i}}$ represents the warped ray due to the camera movement. By setting the number of $\mathcal{N}$ to be less than $K^{2}$, more efficient training is feasible. As the warped rays are determined by parameterized learning, the blurring kernel is deformable, not in a fixed grid manner.
3.2 Continuous Motion Blur Kernel
Continuous Latent Space Modeling for Camera Motion.
Camera-motion blur in images arises from camera movement during exposure time. Such movement, often due to the motion of the hand holding the camera, allows for the representation of camera pose changes as a temporally continuous function, as shown in Fig. 1. Therefore, our goal is to model the continuous movement of the camera. To assign this movement with continuity, we transform the given information of the rays into latent features and apply them to a Neural-ODE [7] to design a continuous latent space.
We embed information about the initial ray of camera motion into the latent space using a parameterized encoder $\mathcal{E}_{\theta}$, utilizing the scene’s view index $idx_{view}$ and 2D pixel location $\mathbf{p}$:
$\mathbf{z}(t_{0})=\mathcal{E}_{\theta}\left(\psi_{v}(idx_{view}),~{}\psi_{p}(\mathbf{p})\right),$ (6)
where $\mathbf{z}_{t_{0}}\in\mathbb{R}^{d}$ is the latent feature of initial ray with hidden dimension $d$, and $\psi_{v}$ and $\psi_{p}$ denote embedding functions for view and pixel information, respectively. Within a small step limit of the latent feature $\mathbf{z}(t)$, the local continuous dynamic is expressed as $\mathbf{z}(t+\epsilon)=\mathbf{z}(t)+\epsilon\cdot d\mathbf{z}/dt$. To apply continuous dynamics to $\mathbf{z}(t)$, we model a parameterized neural derivative function $f$ in the latent space. The derivative of the continuous function is defined as follows:
$\frac{d\mathbf{z}(t)}{dt}=f(\mathbf{z}(t),t;\phi),$ (7)
where $\phi$ denotes the learnable parameters of $f$. With $\mathbf{z}(t_{0})$ and the derivative function $f$, we define an initial value problem (IVP), and the features of subsequential rays in the latent space are obtained by the integral of $f$ from $t_{0}$ to the desired time. This dynamic leads to the format of ODE, and the process of obtaining latent features of the camera motion trajectory using various ODE solvers [12, 24, 11] is expressed as follows:
[…]
where $0\leq n<\mathcal{N}$, and $\mathcal{N}$ is the number of rays in the camera motion. As a result of this approach, we obtain the latent features $\mathbf{Z}\in\mathbb{R}^{\mathcal{N}\times d}$ for $\mathcal{N}$ rays.
Motion-blurred images all have different exposure times, but the given images lack information about them. Assuming a common exposure time for all images when obtaining latent features through the ODE solver may lead to suboptimal local minima since all the images exhibit varying degrees of blur due to different exposure times. Therefore, inspired by the difference in exposure time for images of a single scene, we define a chrono-view embedding function $\mathbf{\Psi}$ that simultaneously embeds given time $t$ and view index $idx_{view}$. Then, $f$ is expressed as follows:
$f(\mathbf{z}(t),t;\phi)\rightarrow f(\mathbf{z}(t),t,\mathbf{\Psi}(t;idx_{view});\phi).$ (9)
Exploiting $\mathbf{\Psi}$, when the conditions of $t_{n}$ and $idx_{view}$ are given to the ODE solver for the IVP, the subsequential latent feature is defined as an unique solution by Picard’s existence theorem [29].
Motion-Blurring Kernel Generation.
The latent features $\mathbf{Z}$ of all rays on the camera motion trajectory, obtained through continuous dynamics modeling, must be decoded into changes in camera pose. We define a decoder $\mathcal{D}_{\xi}$, represented by a shallow MLP parameterized by $\xi$, which outputs three components: the change in 2D kernel location $\Delta\mathbf{p}$, the change in ray origin $\Delta\mathbf{o}_{\mathbf{p}}$, and the corresponding weight $w_{p}$ as defined in Eq. 5:
$(\Delta\mathbf{p}^{t},\Delta\mathbf{o}_{\mathbf{p}^{t}},w^{t}_{p})=\mathcal{D}_{\xi}(\mathbf{z}(t)).$ (10)
Then, $t$-th warped ray $\mathbf{r}^{t}$ is generated from the initial ray $\mathbf{r}=\mathbf{o}+\tau\mathbf{d}$ by following equation:
$\mathbf{r}^{t}=(\mathbf{o}+\Delta{\mathbf{o}_{\mathbf{p}^{t}}})+\tau\mathbf{d}_{\mathbf{p}^{t}},~{}\mathbf{p}^{t}=\mathbf{p}+\Delta\mathbf{p}^{t},$ (11)
where $\mathbf{o}$ and $\mathbf{p}$ denote the ray origin and the 2D pixel location of initial ray, respectively; $\mathbf{d}_{\mathbf{p}^{t}}$ stands for the warped direction by $\mathbf{p}^{t}$. After acquiring sharp pixel colors $\mathbf{c}_{p}(\mathbf{r}^{t})$ for all $\mathcal{N}$ warped rays with inherent continuous camera movement, the blurry pixel color is computed using Eq. 5.
Residual Momentum.
Latent features of the camera motion trajectory are predicted by CMBK and are expected to be optimized sequentially. In this process, predicted origin and direction of the ray $\mathbf{r}^{t}$ may diverge from the initial ray $\mathbf{r}$, potentially leading to a suboptimal kernel. Ma et al. [31] prevents ray divergence by applying a hyperbolic tangent function to $\Delta\mathbf{p}^{t}$ for regularization. However, such regularization is inefficient due to the varying intensity of motion blur across all the images. To address this issue, we apply a residual term to the latent derivative function $f$, implementing regularization that ensures the predicted ray $\mathbf{r}^{t_{i}}$ does not significantly deviate from the previous ray $\mathbf{r}^{t_{i-1}}$:
$f_{\phi}(\mathbf{z}(t_{i-1}))\rightarrow\Lambda(f_{\phi}(\mathbf{z}(t_{i-1}))+\mathbf{z}(t_{i-1})),$ (12)
where the $\Lambda$ is a shallow MLP for regularization and $f_{\phi}(\mathbf{z})$ is the simplified version of $f(\mathbf{z},t,\mathbf{\Psi}(t;idx_{view});\phi)$. This approach prevents the divergence of the camera trajectory, allowing rays to be warped regardless of the motion blur intensity of the image. Note that while the proposed residual momentum is implemented similarly to the methodology of residual connections in ResNet [15], the underlying conceptual frameworks of the two approaches are distinct.
Output Suppression Loss.
We encode the ray and the view information into the latent space using $\mathcal{E}_{\theta}$, and decode it into changes of the ray in camera motion trajectory using $\mathcal{D}_{\xi}$. In this process, since latent feature the initial ray $\mathbf{z}(t_{0})$ serves as the initial value for the ODE, there should be no change in the initial ray. Therefore, we apply an output suppression loss as follows:
$\mathcal{L}_{supp}=\lambda_{supp}\left\|\mathcal{D}_{\xi}(\mathbf{z}(t_{0}))\right\|_{2},$ (13)
where $\lambda$ is the weight for the loss $\mathcal{L}_{supp}$. In this process, the color weight $w_{p}^{t_{0}}$ is not included for the loss. This concept is similar to the cycle consistency in CycleGAN [62]. Minimizing the changes of initial ray to zero value, we ensure that not just the initial ray but also the warped rays do not diverge, providing a regularization effect.
3.3 Optimization
Following Sec. 3.1, our goal is to minimize the difference between the blurry pixel color $\hat{\mathbf{c}}_{blur}(\mathbf{r})$ obtained by the blurring kernel and the ground truth pixel color of the motion-blurred image $\mathbf{c}_{blur}(\mathbf{r})$. Hence, the photometric loss $\mathcal{L}_{photo}$ is represented by $\mathcal{L}_{photo}=\left\|\mathbf{c}_{blur}(\mathbf{r})-\hat{\mathbf{c}}_{blur}(\mathbf{r})\right\|_{2}^{2}$. Furthermore, the parameterized 3D voxel grids $\mathcal{G}_{\sigma}$ and $\mathcal{G}_{c}$ are regularized through the total variation [1] losses $\mathcal{L}_{TV}^{\sigma}$ and $\mathcal{L}_{TV}^{c}$, respectively. Additionally, to prevent divergence of the warped rays and ensure initial ray consistency, Eq. 13 is applied, resulting in our combined objective as follows:
$\mathcal{L}=\mathcal{L}_{photo}+\lambda_{TV}^{\sigma}\mathcal{L}_{TV}^{\sigma}+\lambda_{TV}^{c}\mathcal{L}_{TV}^{c}+\lambda_{supp}\mathcal{L}_{supp}$ (14)
where the $\lambda_{TV}^{\sigma}$ and the $\lambda_{TV}^{c}$ are the weights for $\mathcal{L}_{TV}^{\sigma}$ and $\mathcal{L}_{TV}^{c}$, respectively.
4 Experiments
4.1 Implementation Details
Datasets.
In this experiment, we utilize the camera motion blur dataset published by [31], which is divided into synthetic and real-world scenes. The synthetic scene dataset comprises five scenes synthesized using Blender [8], with multi-view cameras set up to render images with applied camera motion. The camera motion trajectory is formed by linearly interpolating between the poses of the first and last cameras, and the rendered multi-view images are combined in the linear RGB space to obtain the final blurry images. The real-world scene dataset consists of ten scenes captured with an actual camera. The blurry images are obtained by physically shaking the camera during the exposure time, and the camera poses are calculated using the images obtained with COLMAP [46, 45].
Methods Synthetic Scene Dataset Real-World Scene Dataset  Training Time (h)  Rendering Time (s)
 PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
Naive NeRF [33] 23.78 0.6807 0.3362 22.69 0.6347 0.3687 $\sim$ 6 $\sim$ 7
NeRF+MPR [60] 25.11 0.7476 0.2148 23.38 0.6655 0.3140 $\sim$ 6 $\sim$ 7
NeRF+PVD [47] 24.58 0.7190 0.2475 23.10 0.6389 0.3425 $\sim$ 6 $\sim$ 7
Deblur-NeRF [31] 28.77 0.8593 0.1400 25.63 0.7645 0.1820 $\sim$ 25 $\sim$ 7
PDRF-10* [40] 28.86 0.8795 0.1139 25.90 0.7734 0.1825 2.9 0.91
BAD-NeRF* [56] 27.32 0.8178 0.1127 22.82 0.6315 0.2887 $\sim$ 20 $\sim$ 7
DP-NeRF [25] 29.23 0.8674 0.1184 25.91 0.7751 0.1602 $\sim$ 26 $\sim$ 7
SMURF 30.98 0.9147 0.0609 26.52 0.7986 0.1013 1.7 0.77
Table 1: Quantitative comparisons on synthetic and real-world scene dataset. We evaluate the quantitative performance using three metrics, demonstrating that our proposed model significantly ourperforms existing ones even with faster training and rendering times. The red and orange cells respectively indicate the highest and second-highest value.
Figure 4: Qualitative comparisons on synthetic scenes and real-world scenes. Without long training times and sharp images, SMURF produces results most similar to the reference images. SMURF models the detailed aspects that are not captured by previous methods with long training time.
Training Details.
We implement our model on TensoRF [6], an explicit grid-based method, as our backbone renderer. We upsample the voxel counts of grids $\mathcal{G}_{\sigma}$ and $\mathcal{G}_{c}$ from $128^{3}$ to $480^{3}$, and set the feature component $F$ of the grids to 36 and 96, respectively. To implement the CMBK, we set the number of warped rays $\mathcal{N}$ on the camera motion trajectory to 8 and adopt the Euler method [12] as the solver for the Neural-ODE. Single scene training is conducted for 40k iterations, with the learning rates for the grids decaying from $0.02$ to $0.002$, and the learning rate for CMBK decaying from $0.001$ to $0.0001$. The weights for the losses $\lambda_{TV}^{\sigma}$, $\lambda_{TV}^{c}$, and $\lambda_{supp}$ are all set to 0.1. All our experiments are conducted on a single NVIDIA RTX 3090.
Evaluation Metrics.
The images rendered through the SMURF are quantitatively evaluated using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS).
Synthetic Scene Factory CozyRoom Pool Tanabata Trolley
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 19.32 0.4563 0.5304 25.66 0.7941 0.2288 30.45 0.8354 0.1932 22.22 0.6807 0.3653 21.25 0.6370 0.3633
MPR+NeRF 21.70 0.6153 0.3094 27.88 0.8502 0.1153 30.64 0.8385 0.1641 22.71 0.7199 0.2509 22.64 0.7141 0.2344
PVD+NeRF 20.33 0.5386 0.3667 27.74 0.8296 0.1451 27.56 0.7626 0.2148 23.44 0.7293 0.2542 23.81 0.7351 0.2567
Deblur-NeRF 25.60 0.7750 0.2687 32.08 0.9261 0.0477 31.61 0.8682 0.1246 27.11 0.8640 0.1228 27.45 0.8632 0.1363
PDRF-10* 25.87 0.8316 0.1915 31.13 0.9225 0.0439 31.00 0.8583 0.1408 28.01 0.8931 0.1004 28.29 0.8921 0.0931
BAD-NeRF* 24.43 0.7274 0.2134 29.77 0.8864 0.0616 31.51 0.8620 0.0802 25.32 0.8081 0.1077 25.58 0.8049 0.1008
DP-NeRF 25.91 0.7787 0.2494 32.65 0.9317 0.0355 31.96 0.8768 0.0908 27.61 0.8748 0.1033 28.03 0.8752 0.1129
SMURF 29.87 0.8958 0.1057 32.48 0.9285 0.0379 32.34 0.8884 0.0779 29.91 0.9300 0.0436 30.30 0.9307 0.0397
Table 2: Comparison of performance for individual scenes on the synthetic dataset. SMURF exhibits higher performance across all scenes, with the exception of “CozyRoom,” where it shows slightly lower performance relative to others.
Real-World Scene Ball Basket Buick Coffee Decoration
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 24.08 0.6237 0.3992 23.72 0.7086 0.3223 21.59 0.6325 0.3502 26.48 0.8064 0.2896 22.39 0.6609 0.3633
Deblur-NeRF 27.36 0.7656 0.2230 27.67 0.8449 0.1481 24.77 0.7700 0.1752 30.93 0.8981 0.1244 24.19 0.7707 0.1862
PDRF-10* 27.37 0.7642 0.2093 28.36 0.8736 0.1179 25.73 0.7916 0.1582 31.79 0.9002 0.1133 23.55 0.7508 0.2145
BAD-NeRF* 21.33 0.5096 0.4692 26.44 0.8080 0.1325 21.63 0.6429 0.2593 28.98 0.8369 0.1956 22.13 0.6316 0.2894
DP-NeRF 27.20 0.7652 0.2088 27.74 0.8455 0.1294 25.70 0.7922 0.1405 31.19 0.9049 0.1002 24.31 0.7811 0.1639
SMURF 27.50 0.7760 0.1298 28.95 0.8842 0.0619 27.10 0.8409 0.0839 31.33 0.8879 0.0874 24.90 0.8114 0.1190
Real-World Scene Girl Heron Parterre Puppet Stair
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
Naive NeRF 20.07 0.7075 0.3196 20.50 0.5217 0.4129 23.14 0.6201 0.4046 22.09 0.6093 0.3389 22.87 0.4561 0.4868
Deblur-NeRF 22.27 0.7976 0.1687 22.63 0.6874 0.2099 25.82 0.7597 0.2161 25.24 0.7510 0.1577 25.39 0.6296 0.2102
PDRF-10* 24.12 0.8328 0.1679 22.53 0.6880 0.2358 25.36 0.7601 0.2263 25.02 0.7496 0.1532 25.20 0.6235 0.2288
BAD-NeRF* 18.10 0.5652 0.3933 22.18 0.6479 0.2226 23.44 0.6243 0.3151 22.48 0.6249 0.2762 21.52 0.4237 0.3341
DP-NeRF 23.33 0.8139 0.1498 22.88 0.6930 0.1914 25.86 0.7665 0.1900 25.25 0.7536 0.1505 25.59 0.6349 0.1772
SMURF 25.66 0.8592 0.0829 23.59 0.7317 0.1381 25.47 0.7825 0.1207 25.19 0.7702 0.1077 25.48 0.6421 0.0822
Table 3: Comparison of performance for individual scenes on the real-world dataset. While SMURF shows slightly lower PSNR and SSIM in some scenes, it shows significantly higher LPIPS across all scenes compared to previous methods.
4.2 Novel View Synthesis Results
Quantitative Results.
We show the quantitative evaluation of our network, SMURF, in comparison to various baselines on the two benchmark datasets proposed by Deblur-NeRF [31], as shown in Tab. 1. NeRF+MPR and NeRF+PVD are trained with a naive NeRF, where the single-image deblurring methods MPR [60] and PVD [47] to the input data. Additionally, since PDRF-10 [40] does not specify LPIPS in their experiment, we obtain performance of three evaluation metrics using their released code. Furthermore, as BAD-NeRF [56] specifies results from experiments on a modified benchmark dataset, for a fair evaluation, we apply the released code of BAD-NeRF to the benchmark dataset. According to  Tab. 1, despite significantly reduced training and rendering times compared to previous methods, we demonstrate superior quantitative performance across all the metrics. Especially, LPIPS of SMURF demonstrates SMURF’s exceptional perceptual quality across all datasets. Tab. 2 and Tab. 3 detail the individual performance of scenes within the synthetic and real-world datasets, respectively. For synthetic scenes, except for “Cozyroom,” all scenes show quantitatively high performance, with this scene also displaying no significant difference when compared to DP-NeRF. Additionally, for the real-world scenes, despite a few scenes exhibiting somewhat lower PSNR, they demonstrate better perceptual quality through superior LPIPS scores.
Qualitative Results.
We validate the effectiveness of SMURF’s quantitatively high performance through qualitative evaluation via novel view rendering. According to Fig. 4, we compare our results with previous 3D scene deblurring methods for two synthetic scenes (“Factory” and “Tanabata”) and three real-world scenes (“Girl”, “Buick”, and “Basket”). For the “Factory” scene, when comparing the perceptual quality of the reconstructed lowest part of the stairs, it is evident that BAD-NeRF [56] and SMURF best estimate the 3D geometry. However, BAD-NeRF shows an inability to capture the overall color as accurately as other methods. In the rendering results for the “Tanabata” scene, while other methods fail to accurately estimate the position of the power lines, SMURF restores them most similarly to the reference image. In the “Girl” and “Buick” scenes, our method notably restores the visual quality of the shelf’s black support rods and the car logo, respectively, most closely to the reference images. For the “Basket” scene, the holes of the basket further demonstrate the superior qualitative outcome of our results. For additional novel view rendering results, refer to our Appendix.
4.3 Ablation Study
Chrono-View Embedding Function.
To demonstrate the effectiveness of the chrono-view embedding function $\mathbf{\Psi}$ in making continuous latent space modeling robust to exposure time, as discussed in Sec. 3, we conduct various ablative experiments. We conduct experiments by dividing the embedding types into three categories. The time embedding assumes that all images have the same exposure time without incorporating view information, while the chrono-time embedding applies view information, ensuring that all images have distinct exposure time information. As shown in Tab. 4, the application of time embedding yields higher performance in general, but chrono-time embedding shows higher performance only in real-world scenes, while its effect in synthetic scenes is minimal. This is attributed to the characteristics of the synthetic dataset created by Deblur-NeRF [31] using Blender [8], where motion blur images are acquired by linearly interpolating between camera poses to synthesize a total of 10 images. In other words, the synthetic scenes set a constant speed for camera motion throughout the exposure time for all images, implying that all input images share the same exposure time information. Our chrono-view embedding function assumes that all images have varying exposure times, leading to its minimal effect on the synthetic dataset. However, the real-world dataset consists of images captured with an actual camera, where exposure times are likely to vary unless shutter speed is fixed, and the speed of camera motion is not constant during the exposure time. Therefore, applying this function to the real-world scenes yield improved performance due to these variations.
Methods  Embedding Synthetic Scene Dataset Real-World Scene Dataset
Type  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
SMURF None 30.54 0.8995 0.0709 25.74 0.7755 0.1194
SMURF Time 30.94 0.9158 0.0610 26.34 0.7881 0.1102
SMURF Chrono-View 30.98 0.9147 0.0609 26.52 0.7986 0.1013
Table 4: Ablations on chrono-view embedding function. We conduct an ablation study by dividing the embedding types into three categories: no embedding, time embedding, and chrono-time embedding.
Regularzation Strategies.
To validate the effectiveness of the regularization strategies presented in Sec. 3, namely output suppression loss and residual momentum, we conduct various experiments as shown in Tab. 5. The output suppression loss is designed to suppress changes in the origin and direction of the initial ray on the camera motion trajectory to zero, yielding higher performance when applied. This is because the changes in rays are prevented from diverging since the estimated poses also share the same derivative function $f$, similar to the effect of the alignment loss proposed by Deblur-NeRF [31]. This similarity suggests that aligning ray origins is robust to inaccuracies in camera pose, which can be ambiguous due to motion blur. Consequently, Tab. 5 demonstrates that the performance of PSNR and SSIM in synthetic and real-world scenes improves more significantly.
Methods  O.S. Loss Residual Synthetic Scene Dataset Real-World Scene Dataset
 Momentum  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
SMURF 26.17 0.8186 0.0759 25.27 0.7576 0.1121
SMURF 28.49 0.8727 0.0675 25.60 0.7748 0.1057
SMURF 30.72 0.9052 0.0684 25.99 0.7846 0.1181
SMURF 30.98 0.9147 0.0609 26.52 0.7986 0.1013
Table 5: Ablations on regularization strategies. We conduct ablative experiments on the two proposed normalization methods. “O.S. Loss” refers to the output suppression loss.
Residual momentum offers a similar effect to output suppression loss. While the change in camera poses is estimated through a parameterized differential equation, the shape of the derivative obtained in an unsupervised manner could diverge. Therefore, we use residual momentum to ensure the next pose does not deviate significantly from the previous one. Tab. 5 shows that applying another regularization strategy, residual momentum, shows better performance in terms of LPIPS than output suppression loss. Combining two strategies, the SMURF results in best performance across all metrics.
5 Conclusion
We have proposed SMURF, a novel approach that sequentially models accurate camera motion for reconstructing sharp 3D scenes from motion-blurred images. Unlike previous approaches that estimate camera motion in a single step, SMURF incorporates, for the first time, a kernel for estimating sequential camera motions, named the CMBK. This camera motion is represented with continuity by solving continuous dynamics in the latent space using Neural-ODEs. To prevent the divergence of rays estimated by CMBK beyond the motion blur range, we apply regularization techniques: residual momentum and output suppression loss. Furthermore, we model the 3D scene using tensor factorization-based representation, which allows for the integration of incomplete blurred information and complete sharp information within adjacent voxels via CMBK, thereby reducing the uncertainty of blurry information. SMURF significantly outperform previous works quantitatively with faster training and rendering, and its qualitative evaluation is demonstrated through novel view rendering results.
Appendix
Appendix A Details of CMBK
Figure 5: Detailed process for generating kernel. To highlight the continuous dynamics in the latent space, all embedding functions are omitted from the figure. $\mathbf{p}(x_{t_{0}})$ and $\mathbf{p}(y_{t_{0}})$ correspond to the $x$ and $y$ coordinates of the pixel associated with the initial ray, respectively.
We elaborate on the continuous dynamics of the proposed CMBK as shown in Fig. 5. We assume that camera motion encompasses inherent dynamics with a unique solution. The assumed solution is represented by the lighter circles in the left plots of Fig. 5 (b). Rather than implementing the inherent dynamics in a simple physical space, we refine them within a latent space with parametric learning. The continuous dynamics of CMBK involve transforming the pixel coordinates corresponding to the ray into latent features via a parameterized encoder $\mathcal{E}_{\theta}$, and a unique numerical solution [29] is obtained by solving the initial value problem on the formed latent space through a neural ordinary differential equation [7]. The solution obtained represents the change in pixel of the warped ray relative to the initial ray’s pixel. These changes define the pixels corresponding to the warped rays, and we specify Eq. (11) from the main paper:
$\begin{array}[]{ccc}\vspace{+1mm}\mathbf{p}^{t_{i}}=\left(\mathbf{p}(x_{t_{i}}),~{}\mathbf{p}(y_{t_{i}})\right),\\ \vspace{+1mm}\mathbf{p}(x_{t_{i+1}})=\mathbf{p}(x_{t_{i}})+\Delta\mathbf{p}(x_{t_{i+1}}),\\ \mathbf{p}(y_{t_{i+1}})=\mathbf{p}(y_{t_{i}})+\Delta\mathbf{p}(y_{t_{i+1}}).\end{array}$ (15)
Following this principle, the change in the initial ray must necessarily be zero, and to ensure this, the proposed method is the output suppression loss.
Appendix B Number of Warped Rays
Methods  $\mathcal{N}$ Synthetic Scene Dataset Real-World Scene Dataset  Training Time (h)
 PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)  PSNR($\uparrow$)  SSIM($\uparrow$)  LPIPS($\downarrow$)
SMURF 4 29.52 0.8829 0.1014 25.81 0.7710 0.1388 1.16
SMURF 5 30.30 0.9013 0.0926 25.91 0.7811 0.1253 1.27
SMURF 6 29.85 0.8929 0.0851 25.98 0.7822 0.1155 1.43
SMURF 7 30.23 0.8964 0.0751 26.40 0.7944 0.1052 1.56
SMURF 8 30.98 0.9147 0.0609 26.52 0.7986 0.1013 1.72
SMURF 9 30.41 0.9086 0.0575 26.24 0.7922 0.1021 1.88
Table 6: Performance and training time of SMURF according to the number of warped rays. The red, orange, and yellow cells respectively indicate the highest, second-highest, and third-highest value.
We conduct extensive experiments to analyze the performance of the proposed CMBK based on the number of warped rays, $\mathcal{N}$. As shown in Tab. 6, larger $\mathcal{N}$ requires the more pixels to be rendered, resulting in an almost linear increase in training time. Moreover, Fig. 6 shows the performance according to the number of warped rays. Across all datasets, an increase in the number of warped rays tends to improve the PSNR and SSIM metrics. LPIPS noticeably decreases with more rays, interpreting that a higher number of warped rays ensures better perceptual quality. The performance for individual scenes in the synthetic dataset is shown in Tab. 7, showing that except for the “Cozyroom” scene, LPIPS decreases and overall performance peaks when $\mathcal{N}$ is 9. Analysis for the “Cozyroom” scene is conducted in Appendix D. As indicated in Tab. 8, for the real-world dataset, a larger $\mathcal{N}$ generally guarantees higher performance across most scenes. However, there are scenes where performance drops when $\mathcal{N}$ exceeds 9, suggesting that the optimal $\mathcal{N}$ might be 9. Even with $\mathcal{N}$ set to 5, which is the same condition to DP-NeRF [25] and Deblur-NeRF [31], SMURF outperforms them across all metrics. Furthermore, SMURF achieves higher performance with fewer warped rays than PDRF-10 [40], which set $\mathcal{N}$ at 10, validating the effectiveness of our proposed ideas.
Figure 6: Variation in performance with the number of warped rays.
Synthetic  $\mathcal{N}$ Factory CozyRoom Pool Tanabata Trolley
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 25.13 0.7697 0.2127 32.46 0.9294 0.0407 33.04 0.8993 0.0825 28.38 0.9059 0.0825 28.57 0.9101 0.0890
SMURF 5 27.33 0.8381 0.2075 32.72 0.9315 0.0375 32.47 0.8904 0.0797 29.34 0.9222 0.0631 29.65 0.9241 0.0755
SMURF 6 26.74 0.8159 0.1897 31.74 0.9246 0.0381 32.56 0.8922 0.0778 29.22 0.9179 0.0604 28.99 0.9139 0.0599
SMURF 7 26.15 0.8010 0.1679 32.12 0.9269 0.0378 32.59 0.8926 0.0773 29.69 0.9269 0.0490 30.58 0.9348 0.0438
SMURF 8 29.87 0.8958 0.1057 32.48 0.9285 0.0379 32.34 0.8884 0.0779 29.91 0.9300 0.0436 30.30 0.9307 0.0397
SMURF 9 30.52 0.9065 0.0807 31.43 0.9198 0.0428 32.10 0.8849 0.0771 28.93 0.9184 0.0448 29.05 0.9136 0.0423
Table 7: Performance and training time of SMURF for individual scenes in the synthetic dataset, according to the number of warped rays.
Real-World  $\mathcal{N}$ Ball Basket Buick Coffee Decoration
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 26.47 0.7424 0.1754 28.30 0.8653 0.0847 26.57 0.8270 0.1066 30.66 0.8676 0.1228 24.59 0.7972 0.1471
SMURF 5 26.63 0.7528 0.1549 27.96 0.8648 0.0741 26.03 0.8254 0.0947 30.93 0.8734 0.1172 24.22 0.7799 0.1497
SMURF 6 27.31 0.7678 0.1434 27.19 0.8463 0.0773 26.75 0.8315 0.0905 30.76 0.8731 0.1004 24.16 0.7744 0.1465
SMURF 7 27.68 0.7821 0.1325 29.24 0.8862 0.0617 27.05 0.8395 0.0867 30.67 0.8649 0.0931 24.97 0.8092 0.1220
SMURF 8 27.50 0.7760 0.1298 28.95 0.8842 0.0619 27.10 0.8409 0.0839 31.33 0.8879 0.0874 24.90 0.8114 0.1190
SMURF 9 27.16 0.7698 0.1315 28.52 0.8766 0.0631 26.92 0.8366 0.0813 31.41 0.8802 0.0870 24.12 0.7753 0.1405
Real-World  $\mathcal{N}$ Girl Heron Parterre Puppet Stair
 PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS  PSNR  SSIM  LPIPS
SMURF 4 24.66 0.8410 0.1084 23.36 0.7144 0.1863 25.43 0.7644 0.1611 24.63 0.7496 0.1277 23.38 0.5406 0.1687
SMURF 5 24.73 0.8354 0.1028 23.47 0.7244 0.1717 25.22 0.7640 0.1474 24.69 0.7514 0.1239 25.22 0.6391 0.1171
SMURF 6 25.10 0.8473 0.0940 23.60 0.7386 0.1533 24.99 0.7569 0.1402 24.51 0.7493 0.1145 25.38 0.6371 0.0956
SMURF 7 25.43 0.8570 0.0884 23.66 0.7352 0.1457 25.22 0.7744 0.1302 24.70 0.7587 0.1076 25.42 0.6370 0.0846
SMURF 8 25.66 0.8592 0.0829 23.59 0.7317 0.1381 25.47 0.7825 0.1207 25.19 0.7702 0.1077 25.48 0.6421 0.0822
SMURF 9 25.56 0.8567 0.0822 23.81 0.7828 0.1333 24.91 0.7583 0.1216 24.71 0.7554 0.1035 25.31 0.6302 0.0777
Table 8: Performance and training time of SMURF for individual scenes in the real-world dataset, according to the number of warped rays.
Appendix C Additional Rendering Results
Additional rendering results are shown in Fig. 8 (for the synthetic dataset) and Figs. 9 and 10 (for the real-world dataset), demonstrating that our SMURF qualitatively offers the best quality when compared to the reference images.
Appendix D Analysis for Low-PSNR Scenes
Figure 7: Qualitative comparison of rendering results and error maps for the Cozyroom scene. DP-TensoRF is a model that applies the kernel proposed by DP-NeRF [25] to TensoRF [6].
In this section, we analyze individual scenes that showed slightly lower performance from the main paper . Notably, we present the error maps for the “Cozyroom” scene from the synthetic dataset for DP-NeRF [25], DP-TensoRF [25, 6], and SMURF in Fig. 7, where DP-TensoRF is a model that applies the blurring kernel proposed by DP-NeRF to TensoRF. The error map for DP-NeRF shows bright wall rendered cleanly without noise, whereas DP-TensoRF and SMURF exhibit noise on the wall. This indicates that the cause of noise is not the proposed CMBK but the inherent characteristics of the backbone model, TensoRF. As shown in Fig. 8, aside from the noise on the wall, the rendering results of SMURF show slightly sharper image quality in other areas except the wall. Moreover, the “Coffee”, “Parterre”, and “Puppet" scenes from real-world dataset, SMURF shows the best LPIPS score but somewhat lower PSNR. However, when comparing the rendering results in Figs. 9 and 10, it is observable that the results of SMURF are most similar to the reference images for these scenes.
Appendix E Limitations and Future Work
By adopting TensoRF [6], a 3D tensor factorization-based method, as our backbone, we ensure high quantitative performance, superior perceptual quality, and faster training. However, with the advent of 3D Gaussian Splatting [21], which allows for GPU-based rasterization instead of optimizing per ray, backbones that guarantee more faster training and rendering become available. Although our backbone may be slower compared to 3D Gaussian Splatting, applying the main idea of CMBK, continuous dynamics, to a rasterization-based method is expected to result in faster training and rendering. Furthermore, by demonstrating the applicability of continuous dynamics to the 3D scene deblurring, we anticipate the possibility of designing models that cover not only camera motion blur but also object motion blur, which is caused by the movement of objects within the scene.
Figure 8: Qualitative comparison for individual scenes in the synthetic dataset.
Figure 9: Qualitative comparison for individual scenes in the real-world dataset.
Figure 10: Qualitative comparison for individual scenes in the real-world dataset.